Undetermined Coefficients: Find the general solution for the differential equations.
Find the general solution for the following differential equations. (1) y - y – 4y + 4y = 5 - e* + e-* (2) y + 2y + y =
Answer #1

A لا يطي 0 y y _uy+uy s-e²+ e* -6 Mexitary equatiom m?.-?- um+420 (m-1) (m2-4) = 0 -7 m = 1,2,- The soletin Corresponding